### v1

# Embodied Question Answering via Multi-LLM Systems

Abstract
Embodied Question Answering (EQA) is an important problem, which involves an agent exploring the environment to answer user queries. In the existing literature, EQA has exclusively been studied in single-agent scenarios, where exploration can be time-consuming and costly. In this work, we consider EQA in a multi-agent framework involving multiple large language models (LLM) based agents independently answering queries about a household environment. To generate one answer for each query, we use the individual responses to train a Central Answer Model (CAM) that aggregates responses for a robust answer. Using CAM, we observe a $50\%$ higher EQA accuracy when compared against aggregation methods for ensemble LLM, such as voting schemes and debates. CAM does not require any form of agent communication, alleviating it from the associated costs. We ablate CAM with various nonlinear (neural network, random forest, decision tree, XGBoost) and linear (logistic regression classifier, SVM) algorithms. Finally, we present a feature importance analysis for CAM via permutation feature importance (PFI), quantifying CAMs reliance on each independent agent and query context.
Embodied Question Answering via Multi-LLM Systems
Bhrij Patel University of Maryland, College Park                        Vishnu Sashank Dorbala University of Maryland, College Park Amrit Singh Bedi University of Central Florida                        Dinesh Manocha University of Maryland, College Park
1 Introduction
Embodied Question Answering (EQA) is an important task where an agent explores an environment, such as a household, to answer questions based on its observations. It has gained considerable interest in the past few years due to its applications for in-home robots and personalized assistants Das et al. (2018); Gordon et al. (2018); Duan et al. (2022a); Sima et al. (2023); Luo et al. (2023); Dorbala et al. (2024a); Zhu et al. (2023). For a seamless interface with humans, robotic agents must be equipped with natural language understanding Xu et al. (2021); Sun et al. (2024) and possess common-sense reasoning of the world. One form of common sense is having a relational understanding of household items and their usual place (i.e., a mug is highly likely to be in the kitchen or living room). Foundation models, specifically Large Language Models (LLMs) such as GPT Achiam et al. (2023) and LLama Touvron et al. (2023), have shown remarkable prowess in both natural language understanding and common sense reasoning. Recently, prior art has used these foundation models in EQA for query generation Dorbala et al. (2024a) and answering (Majumdar et al., 2024).
In existing works, only a single agent explores the environment with a limited field of view, inducing low coverage and high exploration costs. In contrast, for an LLM-based embodied agent to succeed in EQA, efficient exploration of the environment is needed as the questions will center around the state of the environment Das et al. (2018); Yu et al. (2019); Zhu et al. (2023); Majumdar et al. (2024). Prior work has studied the zero-shot capabilities of LLM-based agents to explore an unknown household environment Dorbala et al. (2023, 2024b); Majumdar et al. (2024).
The rise of Embodied Artificial Intelligence (EAI) Duan et al. (2022b) brings up the possibility of co-habited spaces with multiple agents and humans working in conjunction Puig et al. (2023). In this context, we consider a similar co-inhabited setup, and propose using a team comprising of multiple LLM-based agents or a Multi-LLM system for EQA. A Multi-LLM system allows for simultaneous exploration of the household environment in the same amount of time.
A key challenge in a multi-agent setup is with conflicting responses when agents are asked the same query. Since LLMs are blackboxes capable of generating free-form natural language, their responses to the same query may significantly vary Gal and Grosz (2022); Wang et al. (2024). In an embodied setting, such responses are further skewed by partial observations due to limited exploration. Prior work in ensemble LLM methods tackle this issue using weighted sum or voting schemes Li et al. (2024); Wei et al. (2022); Wang et al. (2022); Yang et al. (2024) and consensus-reaching debates between the LLM agents Chen et al. (2023a); Du et al. (2023); Xiong et al. (2023); Guo et al. (2024a); Chan et al. (2023). However, such methods can be vulnerable to scenarios where incorrect answers from malicious or poor-performing agents can influence the overall decision. An ideal EQA model would learn to identify incorrect agents and avoid being influenced by their responses. Furthermore, we would like to ideally avoid the communication costs incurred by debate-based approaches.
Main Results: To address the issues presented above, we solve EQA using a novel Multi-LLM agent approach. We instruct the agents to answer based on their observations a set of binary embodied questions about the household. We then train a Central Answer Model (CAM) on the answers of independent agents with supervised learning. The network directly outputs the final answer to a given question and does not require any communication between agents. Furthermore, our model learns which agents to rely on while producing its final answer, which reduces the end user’s effort in having to individually figure out the best agent response. Our novel contributions include:
  • Novel Multi-LLM EQA Framework: We propose a Central Answer Model (CAM) for Embodied Question Answering in a Multi-Agent setting. CAM is a classifier that aggregates multiple agent responses to predict an answer. We generate queries for various Matterport 3D Chang et al. (2017) environments. We train CAM on the labeled query dataset with multiple machine learning methods and show that our best model can achieve up to 50% higher accuracy than a majority vote and debate aggregation methods.
  • Integration with Exploration Systems: We evaluate our multi-agent framework on data gathered using an LLM-based exploration method Dorbala et al. (2023) on multiple agents in Matterport3D, demonstrating our system can work in conjunction with SOTA exploration methods in unknown environments.
  • Examining Agent Reliance: We present a feature importance analysis of CAM reliance on each independent agent using the metric, permutation feature importance (PFI) Breiman (2001); Fisher et al. (2018).
The rest of the paper is organized as follows: in Section 2, we provide a brief overview of the prior art relating to EQA and Multi-LLM systems; Section 3 gives a detailed overview of our approach; in Section 4, we describe our experimentation detail to evaluate our approach and follow with results showing the performance benefits compared to majority vote and consensus-reaching debate. Finally, in Section 7 and 8, we provide concluding remarks and a discussion on limitations, respectively.
Figure 1: Overview of our Multi-LLM framework. LLM-based agents that have explored the dynamic household are independently asked a binary “Yes/No” question and must answer based on their observations. Only one of the agents has the correct answer (“Yes”) while the other two agents are wrong due to either not exploring the relevant portions of the house enough or being malicious. The central network, however from supervised training, has learned to weigh the decision of the first agent more when given context about the item, room, and timestep.
Reference QA Exploration LLM-based Multi-Agent Feature Importance
(Majumdar et al., 2024)
(Singh et al., 2024)
Our Approach
Table 1: Table of our contributions and novelties compared to recent prior work. This work is the first to tackle Embodied EQA with LLM-based Multi-agent system and to provide a feature importance analysis on the framework.
2 Related Works
2.1 Embodied Question Answering
Embodied Question Answering (EQA) was first introduced by Das et al. (2018) and has become a popular task for robotics (Gordon et al., 2018, 2019; Duan et al., 2022a; Sima et al., 2023; Wijmans et al., 2019; Chen et al., 2019; Luo et al., 2019), where traditionally an agent is given a query such as “was there a plant in the living room?’ and explores its environment, in this case a house, to answer the question. Zhu et al. (2023) had alternated between an exploration phase and then a question phase and repeated the process until the query was answered. In our work, we first have an exploration gathering phase, then answer a set of questions. Further work has extended EQA to scenarios where the final answer revolves around multiple objects Yu et al. (2019), and where a target object is not explicitly stated in the question Tan et al. (2023). Tan et al. (2020) introduced multi-agent exploration and cooperation for question-answering in interactive environments. Dorbala et al. (2024a) provided a dataset of subjective or situational queries and was the first to use a generative approach for query creation, and Majumdar et al. (2024) designed the first open-vocabulary EQA dataset to test and benchmark foundation models on this task. Recently, Singh et al. (2024) studied the zero-shot LLM performance on visual QA with indoor house scenes. However, the LLM is given the scenes without any exploration of the environment. We evaluate our Multi-LLM approach with a SOTA LLM-based exploration method for observation gathering.
2.2 Multi-LLM Systems
Research in Multi-LLM systems have grown in popularity with the rise of foundation models for areas such as task planning Chen et al. (2023b), application development Wu et al. (2023), and chatbots Duan et al. (2023). Extensive studies have focused on their abilities to cooperate. Prior work has used multi-LLM debate to reach a consensus (Chan et al., 2023; Du et al., 2023; Guo et al., 2024a). Chen et al. (2023a) analyzes the consensus-seeking debate between multiple LLM agents with different personalities. For embodied AI, Guo et al. (2024b) have shown that multiple LLM-based agents can cooperate in teams by designating a leader for increased task efficiency. Other work have looked into using voting metrics of independent LLM-agents to arrive at a final answer (Li et al., 2024; Wei et al., 2022; Wang et al., 2022). Our work aims to remove the need for communication by training a central classifier on independent answers from multiple agents. Training a classifier in a supervised manner helps our approach become robust to unreliable agents that may influence overall answers in voting schemes.
Figure 2: Overview of experimentation pipeline. [TOP] We have an LLM answer a collection of embodied questions about the Matterport3D environment using the observations collected and stored by a specific agent during an exploration phase. The answers from the LLM are stored offline. [BOTTOM] In the central network training phase, we select multiple sets of answers from different agents to train the classifier with the ground-truth answer with supervised learning.
3 Our Approach: Multi-LLM System with Central Answer Network
In this work, we implement a two-step procedure that comprises a question-answering phase, and a central model training phase. The question-answering phases is agent-specific step while the classification-training step relies on combining the intermediate results of the question-answering phases from various agents. Before starting the question-answering phase, the agents have explored the environment to gather observations. Later, we see the performance of our approach when using a LLM-based SOTA method for exploration. We describe both phases in detail below.
3.1 Question-Answering Phase
In this phase, we generate a dataset of binary embodied questions specific to a Matterport 3D environment. Each LLM-based agent answers based on its independent observations from its exploration of the environment. To generate these questions we follow this template:
Do you think there is a [ITEM] in the [ROOM]? Use both common-sense reasoning about the object and room and the observation list given. Even if the relevant information is not in your observations, the answer can still be YES. Respond with YES or NO.
We then prompt the LLM to answer each question based on the stored observations of the agent. We pass the observations to the system prompt formatted as a dictionary, {ROOM NAME: [LIST OF ITEMS IN THE ROOM]}. The user prompt is a question following the template above and the system prompt is as follows:
You are an embodied agent that has explored a house. I prefer definite answers to questions which might require to explore. The observations are: [OBSERVATION DICTIONARY]
3.2 Central Model Training Phase
To give the final answer to each question, we train a model to give a final “Yes/No" output. The model input dimension is $2+|S|$, and the first two inputs account for the vectorized question in the form of “[object, room]”. The rest of the $|S|$ inputs come from entering the independent answers from each of the agents chosen for the multi-LLM system. During inference, we input unseen questions and agent answers into the network to output a test-time answer. Due to the modularity of our approach, for each set of independent responses, we train multiple machine learning models to produce multiple central answer models to compare against baseline aggregation approaches.
4 Experimentation
We now describe our experimentation in detail. With these experiments, we aim to 1) measure the accuracy of CAM for EQA and show its performance over other aggregation baselines (Section 4.2.2; 2) execute our Multi-LLM framework with observations gathered from a SOTA exploration method for unknown household environments, demonstrating its practicality for real-world scenarios (Section 5); and 3) provide a feature importance analysis to highlight the benefit of utilizing independent agent responses in our approach (Section 6) In all experiments we use GPT-4-turbo.
4.1 Query Generation
We extract the embodied question bank from the Matterport3D environments. For each scan, we extract an offline list of observed items detected in each image of the environment using the GLIP object detection algorithm Li et al. (2022).
For each item and room combination, we formulate the question using the template described in Section 3.1. Because these questions were extracted based on the observed items in the scan, we know the ground truth answer is “Yes".
To create questions with ground-truth answers “No", for each query generated with a “Yes" ground-truth, we formulate a question with the same item but with a random room that is not within the subset of rooms that contains that specific item.
4.2 EQA Performance Evaluation
In this section, to focus on the aggregation abilities of CAM, we have an oracle exploration where we explicitly say which rooms and items an agent has in its observations memory when answering the questions. In Section 5 we use a SOTA algorithm, LGX Dorbala et al. (2023) for the exploration phase.
Metric. To measure and compare the performances of our CAM approach against baseline aggregation approaches, we calculate the inference time accuracy, or correctness, of the model:
$\frac{1}{N}\sum_{i=1}^{N}\mathbbm{1}_{y_{i}=\hat{y}_{i}},$ (1)
where $y_{i}$ is the ground truth label of the test question, and $\hat{y}_{i}$ is the final answer prediction from CAM. We produce CAM models with various machine learning algorithms described in Table 2.
Algorithm Linearity
Neural Network (NN) Nonlinear
Random Forest (RF) Nonlinear
Decision Tree (DT) Nonlinear
XGBoost (XG) Nonlinear
RBF SVM (SVM) Nonlinear
Linear SVM (SVM-L) Linear
Logistic Regression (LR) Linear
Table 2: Table of different machine learning algorithms we implement to train CAM. We have a mix of linear and non-linear, parametric and non-parametric methods to thoroughly ablate the central model training phase.
4.2.1 Baselines
We compare these different variations of CAM against two other forms of the aggregation methods used for ensemble LLM methods. Neither method is learning-based and requires not training set. We describe them below:
  • Majority Vote (MV): For each test question, we take the majority vote of the independent answers from the agents of the Multi-LLM system. We do not directly take into account room or object with a majority vote
  • Debate: For each test question, we have the agents of the Multi-LLM system participate in a turn-based discussion where they relay their thoughts to one another for 2 rounds. To initial the debate we provide each LLM-based agent context with the following system prompt.
    System prompt: You are an embodied agent in a house. Your id is [INDEX]. Here are your observations from exploring the house: [OBSERVATION DICTIONARY]. Your initial answer to whether or not there was a [OBJECT] in the [ROOM] was [INITIAL ANSWER]. Other agents may have different answers. Please debate with the other agents to come to a consensus. Here is the conversation history: [CONVERSATION HISTORY]
    The last sentence is replaced with This is the beginning of the conversation. at the start of the debate when no conversation history is present. Below, we provide the user prompt to get the agent’s response for its turn during that round.
    User prompt: It is your turn to speak. Use your observation and conversation history to help.
    When prompted for the final answer at the end of the debate. The user prompt is Please give your final "Yes/No" answer. Give a definite answer.
We then collect the final answers from each agent and take the majority vote as the final prediction.
Figure 3: Outperforms Baselines; Test-time EQA accuracy measured by Equation 1 of CAM methods (blue) and baseline aggregation methods (orange). We report the mean and standard deviation over 5 seeds. The Multi-LLM system in these experiments uses independent answers from three agents that observations from bedroom, dining room, and kitchen. We use a 95%-5% train-test random split. We experiment on Matterport environments with [LEFT] 215 nodes and 15 distinct rooms [RIGHT] 53 nodes and 12 distinct rooms. Against the majority vote (MV) baseline, most CAM methods perform noticeably better during test time, with XGBoost (XG) achieving 50% higher accuracy. Against the debate baseline, the non-linear models perform at least as well, with XGBoost achieving 33% higher accuracy. The range in CAM methods performance highlights the need for careful model selection and tuning and that having ground truth labels does not guarantee significant improvement against the non-learning based baselines.
Figure 4: More Agents, Larger Test Set; With the $215$-node environment, we use a 80%-20% train-test random split, reporting the mean and standard deviation over 5 seeds. The Multi-LLM system in these experiments uses independent answers from $7$ agents. Each agent only used observations from one room to respond to queries. The $7$ rooms used that observations from the bedroom, dining room, living room, office, bathroom, family room and kitchen. Against the majority vote (MV) baseline, all CAM methods perform noticeably better during test time, with XGBoost (XG) achieving 46% higher accuracy.
Figure 5: LGX + CAM; Performance of CAM with $3$ independent agent that performed LGX for collecting observations. We use a 95%-5% train-test random split as in Figure 3. We experiment on Matterport environments with [LEFT] 215 nodes and [RIGHT] 53 nodes. Against both baselines, all CAM methods outperform during test-time inference, with XGBoost having the highest accuracy.
4.2.2 Results
We see the superior performance of CAM in Figure 3, where we run two sets of experiments, one to compare against MV and debating baseline in an $215$-node Matterport environment (left image) and the other in a $53$-node (right) environment. For each experiment, we report the mean and standard deviation accuracy calculated by Equation 1 over $5$ random seeds. We generate $600$ and $480$ queries for the $215$ and $53$-node environments, respectively, using the method described in Section 4.1. For both experiments, we train with $3$ independent answers to each query. The $3$ answers are from there LLM, where each time it answers it has only observations from one of the three rooms: bedroom, dining room, kitchen.
For the Central Model Training phase, the train-test random split of the generated queries and independent LLM-generated answers is $95$%-$5$%. In both environments, majority of the CAM method outperforms the MV and debating baselines. Specifically, XGBoost performs the best with $50$% $33$% higher accuracy than MV and debating, respectively. We also highlight the difference in time costs between the debate baseline and CAM. For each test question where three agents debate for $2$ rounds, the average inference time to make a final prediction is $30$ seconds, while the average time to train and run inference on the test set for XGBoost is $1.3$ seconds.
For a comparison on a larger test set with more input features using the $215$-node environment, in Figure 4 we have a random $80$%-$20$% train-test split. In this experiment, we have $7$ independent answers. Each answer is generated from the LLM, where the LLM only has access to the observations from a single room. The $7$ rooms are bedroom, kitchen, dining room, living room, family room, bathroom, and office. Each CAM method outperforms the baseline MV with XGBoost performing 46% higher.
We note that the two consistently lowest performing CAM methods and the ones closest to the MV and debating baselines are the linear models, Linear SVM and Logistic Regression Classifier. Intuitively, this similarity in performance makes sense as MV can be interpreted as a linear model with fixed equal weights, whereas Linear SVM and LR learn and adjust their weights based on the training data. Since we end up doing a majority vote on the final answers after the discussion, the debate baseline can also be construed as a linear model. These linear methods perform the worst suggesting the non-linearity of this EQA dataset.
5 CAM + Language-Guided Exploration
In this section, we evaluate our Multi-LLM framework by gathering observations using a SOTA exploration method. Rather than manually selecting which observations to provide to the LLM to answer the queries, we have all the LLM-based agents explore the household environment for a set amount of time before answering question.
In the exploration phase, each subordinate agent is set in the Matterport house scan graph and performs an augmented variation of Language-Guided Exploration (LGX) as proposed by Dorbala et al. (2023). In traditional LGX, the LLM agent observes the different immediate household items around itself and chooses which one to navigate towards to find a specific target item, forming a chain of “hops” to the item.
In our variation, we specifically provide the LLM with a system prompt that encourages the agent to explore as much of the house as possible. You are an embodied agent in a house. You want to aggressively explore the house and you want to find as many unique objects as you can.
Query Subject $\downarrow$ Object Kitchen Dining Room Bedroom Office
Kitchen 27.50 22.50 27.50 27.50 22.50
Dining Room 17.50 12.50 12.50 17.50 12.50
Bedroom 25.00 25.00 25.00 25.00 25.00
Office 12.50 5.00 5.00 12.50 7.50
Table 3: This table displays the feature importance of different decision tree-based CAM models. We see in cases where the queries only pertaining to the dining room (third row) that responses based only on dining room observation (fourth column) is not the most important feature. This insight suggests that CAM learned not to rely on that LLM agent.
The agent then travels from node to node in the house graph. For each timestep, we document the timestep, node, and items observed. When the agent reaches a node we can then access the stored results and add them to the list of observed items. We have the $3$ agents perform $10$ steps of this augmented LGX and store the observations. The rest of the experiment process follows the same setup as described in Section 3 with the Question-Answering and Central Model Training phases. We perform both these experiments in the same two environments as in Section 4.2.2 and using a $95\%$-$5\%$ random train-test split over 5 seeds.
The performance of the various CAM methods and baselines in the larger $215$-node graph is shown on the left image of Figure 5 and the performance for the $53$-node environment is shown on the right. From both plots, we again see CAM model outperforms both baselines as with Section 4.2.2. In fact, with LGX for observation collection, we see that on average, all CAM methods have a higher test time accuracy than the non-learning aggregation baselines.
Interestingly, in the 53-node environment debating baselines do slightly worse on average than MV. This degradation occurs due to agents who were initially correct being swayed by incorrect agents during conversation. Overall, our Multi-LLM system can be utilized in conjunction with an LLM-based exploration method, emphasizing its practicality in real-world use cases. In the next section, we provide a feature importance of CAM.
6 Feature Importance Analysis
In the context of EQA, feature importance is a metric by which input impacts the model’s final answer. For our Multi-LLM system, feature importance can be used to examine if any agent response or query context (i.e. room or object) is relied heavily upon for training the central answer model. To examine which features hold the most weight in the final decision of CAM model, we look into studying the variable, or feature importance Breiman (2001); Fisher et al. (2018), of the CAM model. We measure feature importance using permutation feature importance (PFI). We define the PFI the $i$-th feature as,
$PFI(i)=|a_{\text{val}}(m)-a_{\text{val}}^{i}(m)|,$ (2)
where $a_{\text{val}}(m)$ is the model $m$ accuracy measured by Equation 1 of the original validation set, and $a_{\text{val}}^{i}(m)$ is the accuracy with a modified validation set where the $i$-th feature column is randomly permuted. Intuitively, the more important a variable, the more the accuracy should change if the values are randomly shuffled because the variable is highly correlated with the target variable. In our Multi-LLM system for EQA, an agent that gives nearly-deterministic answers (e.g. “No” for almost all answers) would have a PFI close to 00.
To examine feature importance, we train multiple DT-based CAM models, each with a different filtered query dataset from the $215$-node environment. In each training, the features comprise of the room and object name to represent the query and $4$ independent responses from LLMs that had observations either from the kitchen, dining room, bedroom, or office. The dataset however will only consist of rows with a given “Room” value. The model is trained and tested, and we then calculate feature importance using Equation 2.
Table 3 summarizes our feature importance findings. The leftmost column specifies the room the dataset was filtered for, and the other columns show the average PFI of the object feature and independent responses for that filtered dataset over 5 trials. We see that in many cases, like when trained only on the dining room queries, the corresponding feature (i.e. response based only on dining room observation) does not have the highest PFI. We analyzed the offline LLM-generated answers based only on dining room observations and out of the $600$ responses, $549$ are “No”, or $91.5\%$. Thus, the CAM model has learned not to rely on these answers as much as the other responses even though they are based only on observations from the room in question. This outcome demonstrates how analyzing feature importance can help examine individual members of a multi-agent setup and further supports our approach utilizing independent agent responses.
7 Conclusion
In this paper, we tackle Embodied Question Answering with a novel Multi-LLM framework that aggregates the independent responses of LLM-based agents and produces a final answer. To the best of our knowledge, we are the first to use multiple LLM-based agents for EQA. We propose a central answer model (CAM) that learns to weigh the responses of each agent to decide on a single answer to the given query. We show that various machine learning methods are suitable to train CAM and consistently outperform non-learning-based aggregation methods for ensemble outputs, specifically majority voting and multi-LLM debating. A key benefit of our system is that it requires no communication during inference, reducing time costs. Furthermore, our approach can be combined with SOTA methods for exploration in unknown households. We end this work by providing a feature importance analysis to study the effect of individual agents and target object specification on the final answer. We hope that this work will encourage further research in Multi-LLM approaches for EQA as LLM-based embodied agents become more household.
8 Limitations and Future Work
This work has several limitations. First, the ground truth labels to train CAM can be difficult to obtain, especially if the household environment is constantly changing with non-stationary items. We leave it as future work to adapt this Multi-LLM setup for EQA in dynamic household environments. Second, our queries are all objective binary “Yes/No” questions. To better enhance practicality, future work should look into aggregation methods for situational, subjective, and non-binary questions. It would also be useful to remove the need to explicitly mention the target object in the query. Other areas directions for research could be then to apply our CAM aggregation method to question-answering tasks outside of Embodied AI such as long video understanding. Third, our results hinge on the quality of the observation data. In our work, we relied on the GLIP model. Future work should analyze the effect of scene understanding algorithms on our setup.