### v4

# Central Answer Modeling for an Embodied Multi-LLM System

Abstract
Embodied Question Answering (EQA) is an important problem, which involves an agent exploring the environment to answer user queries. In the existing literature, EQA has exclusively been studied in single-agent scenarios, where exploration can be time-consuming and costly. In this work, we consider EQA in a multi-agent framework involving multiple large language models (LLM) based agents independently answering queries about a household environment. To generate one answer for each query, we use the individual responses to train a Central Answer Model (CAM) that aggregates responses for a robust answer. While prior Question Answering (QA) work has used a central module based on answers from multiple LLM-based experts, we specifically look at applying this framework to embodied LLM-based agents that must physically explore the environment first to become experts on their given environment to answer questions. Our work is the first to utilize a central answer model framework with embodied agents that must rely on exploring an unknown environment. We set up a variation of EQA where instead of the agents exploring the environment after the question is asked, the agents first explore the environment for a set amount of time and then answer a set of queries. Using CAM, we observe a $46\%$ higher EQA accuracy when compared against aggregation methods for ensemble LLM, such as voting schemes and debates. CAM does not require any form of agent communication, alleviating it from the associated costs. We ablate CAM with various nonlinear (neural network, random forest, decision tree, XGBoost) and linear (logistic regression classifier, SVM) algorithms. We experiment in various topological graph environments and examine the case where one of the agents is malicious and purposes contributes responses it believes to be wrong. Finally, we present a feature importance analysis for CAM via permutation feature importance (PFI). We provide and code and query dataset for further research.
Introduction
Embodied Question Answering (EQA) is an important task where an agent explores an environment, such as a household, to answer questions based on its observations. It has gained considerable interest in the past few years due to its applications for in-home robots and personalized assistants (Das et al. 2018; Gordon et al. 2018; Duan et al. 2022a; Sima et al. 2023; Luo et al. 2023; Dorbala et al. 2024a; Zhu et al. 2023). For a seamless interface with humans, robotic agents must be equipped with natural language understanding (Xu et al. 2021; Sun et al. 2024) and possess common-sense reasoning of the world. One form of common sense is having a relational understanding of household items and their usual place (i.e., a mug is highly likely to be in the kitchen or living room). Foundation models, specifically Large Language Models (LLMs) such as GPT (Achiam et al. 2023) and LLama (Touvron et al. 2023), have shown remarkable prowess in both natural language understanding and common sense reasoning. Recently, prior art has used these foundation models in EQA for query generation (Dorbala et al. 2024a) and answering (Majumdar et al. 2024).
In existing works, only a single agent explores the environment with a limited field of view, inducing low coverage and high exploration costs. In contrast, for an LLM-based embodied agent to succeed in EQA, efficient exploration of the environment is needed as the questions will center around the state of the environment (Das et al. 2018; Yu et al. 2019; Zhu et al. 2023; Majumdar et al. 2024). Prior work has studied the zero-shot capabilities of LLM-based agents to explore an unknown household environment (Dorbala, Mullen Jr, and Manocha 2023; Dorbala et al. 2024b; Majumdar et al. 2024).
The rise of Embodied Artificial Intelligence (EAI) (Duan et al. 2022b) brings up the possibility of co-habited spaces with multiple agents and humans working in conjunction (Puig et al. 2023). In this context, we consider a similar co-inhabited setup, and propose using a team comprising of multiple LLM-based agents or a Multi-LLM system for EQA. A Multi-LLM system allows for simultaneous exploration of the environment in the same amount of time.
A key challenge in a multi-agent setup is with conflicting responses when agents are asked the same query. Since LLMs are blackboxes capable of generating free-form natural language, their responses to the same query may significantly vary (Gal and Grosz 2022; Wang et al. 2024). In an embodied setting, such responses are further skewed by partial observations due to limited exploration. Prior work in ensemble LLM methods tackle this issue using weighted sum or voting schemes (Li et al. 2024; Wei et al. 2022; Wang et al. 2022; Yang et al. 2024) and consensus-reaching debates between the LLM agents (Chen et al. 2023a; Du et al. 2023; Xiong et al. 2023; Guo et al. 2024a; Chan et al. 2023; Li, Patel, and Du 2023). However, such methods can be vulnerable to scenarios where incorrect answers from malicious or poor-performing agents can influence the overall decision (Amayuelas et al. 2024).
Figure 1: Overview of our Multi-LLM framework. LLM-based agents that have explored the dynamic household are independently asked a binary “Yes/No” question and must answer based on their observations. Only one of the agents has the correct answer (“Yes”) while the other two agents are wrong due to either not exploring the relevant portions of the house enough or being malicious. The central network, however from supervised training, has learned to weigh the decision of the first agent more when given context about the item, room, and timestep.
An ideal EQA model would learn to identify incorrect agents and avoid being influenced by their responses. Furthermore, we would like to ideally avoid the communication costs incurred by debate-based approaches.
Main Results: To address the issues presented above, we apply a Multi-LLM system to solve this embodied task. We instruct the agents to answer based on their observations a set of binary embodied questions about the household. We then train a Central Answer Model (CAM) on the answers of independent agents with supervised learning. The network directly outputs the final answer to a given question and does not require any communication between agents. Furthermore, our model learns which agents to rely on while producing its final answer, which reduces the end user’s effort in having to individually figure out the best agent response. Our novel contributions include:
  • Multi-LLM QA Framework with Embodied Navigation: We train a Central Answer Model (CAM) for Question Answering in an Embodied Multi-Agent setting. We depart from prior EQA setups where the agents explore after being asked a question. Rather we have the agents explore before being asked a set of question on their environment. CAM is a classifier that aggregates multiple agent responses to predict an answer. We generate queries for various Matterport 3D (Chang et al. 2017) environments. We train CAM on the labeled query dataset with multiple machine learning methods and show that our best model can achieve up to 46% higher accuracy than a majority vote and debate aggregation methods. While prior work has used a central answer module to output a final answer from a mixture of answers from multiple LLM experts (Puerto, Şahin, and Gurevych 2023; Si et al. 2023; Jiang, Ren, and Lin 2023), we are the first to apply the framework to embodied agents that must first explore an environment.
  • Integration with Exploration Systems: We evaluate our multi-agent framework on data gathered using an LLM-based exploration method (Dorbala, Mullen Jr, and Manocha 2023) on multiple agents in Matterport3D, demonstrating our system can work in conjunction with SOTA exploration methods in unknown environments.
  • Examining Agent and Feature Reliance: We present a feature importance analysis of CAM reliance on each independent agent using the metric, permutation feature importance (PFI) (Breiman 2001; Fisher, Rudin, and Dominici 2018). We also look at the agreement between the individual agent and the final answer of the central answer model and the baselines. Similarly to Si et al. (2023), some of our methods for the central answer model are interpretable and provide a Decision Tree visualization of one our models.
Reference Exploration Central Module Multi-Agent Feature Importance
(Majumdar et al. 2024)
(Puerto, Şahin, and Gurevych 2023)
Our Approach
Table 1: Table of our contributions and novelties compared to recent prior work. This work is the first to tackle Embodied EQA with LLM-based Multi-agent system and to provide a feature importance analysis on the framework.
Related Works
Embodied Question Answering and Exploration
Embodied Question Answering (EQA) was first introduced by Das et al. (2018) and has become a popular task for robotics (Gordon et al. 2018; Gordon, Fox, and Farhadi 2019; Duan et al. 2022a; Sima et al. 2023; Wijmans et al. 2019; Chen et al. 2019; Luo et al. 2019), where traditionally an agent is given a query such as “was there a plant in the living room?’ and explores its environment, in this case a house, to answer the question. Zhu et al. (2023) had alternated between an exploration phase and then a question phase and repeated the process until the query was answered. Further work has extended EQA to scenarios where the final answer revolves around multiple objects (Yu et al. 2019), and where a target object is not explicitly stated in the question (Tan et al. 2023). Tan et al. (2020) introduced multi-agent exploration and cooperation for question-answering in interactive environments. Dorbala et al. (2024a) provided a dataset of subjective or situational queries and was the first to use a generative approach for query creation, and Majumdar et al. (2024) designed the first open-vocabulary EQA dataset to test and benchmark foundation models on this task. Recently, Singh, Pavlakos, and Stamoulis (2024) studied the zero-shot LLM performance on visual QA with indoor house scenes. However, the LLM is given the scenes without any exploration of the environment. We evaluate our Multi-LLM approach with a SOTA LLM-based exploration method for observation gathering introduced by Dorbala, Mullen Jr, and Manocha (2023).
Note: In our work, we deviate from the traditional EQA approach. We first have an exploration gathering phase, then answer a set of questions on the same environment. We do so for two reasons: 1) to provide a setup to better compare against Multi-LLM QA baselines: majority voting and debating; 2) we believe this setup is better suited for the scenario where the user asks a question about the items in their homes and quickly needs an answer.
Multi-LLM Systems and Experts
Research in Multi-LLM systems have grown in popularity with the rise of foundation models for areas such as task planning (Chen et al. 2023b), application development (Wu et al. 2023), and chatbots (Duan et al. 2023). Extensive studies have focused on their abilities to cooperate. Prior work has used multi-LLM debate to reach a consensus (Chan et al. 2023; Du et al. 2023; Guo et al. 2024a; Chen, Saha, and Bansal 2023). Chen et al. (2023a) analyzes the consensus-seeking debate between multiple LLM agents with different personalities, and Amayuelas et al. (2024) investigates the susceptibility of a multi-LLM discussion to an adversarial agent that purposefully tries to lead an incorrect answer choice. For embodied AI, Guo et al. (2024b) have shown that multiple LLM-based agents can cooperate in teams by designating a leader for increased task efficiency. Other work have looked into using voting metrics of independent LLM-agents to arrive at a final answer (Li et al. 2024; Wei et al. 2022; Wang et al. 2022). Our work aims to remove the need for communication by training a central classifier on independent answers from multiple agents. Training a classifier in a supervised manner helps our approach become robust to unreliable agents that may influence overall answers in voting schemes.
A central architecture to output a final answer from multiple answers from different LLM experts has been explored previously (Puerto, Şahin, and Gurevych 2023; Si et al. 2023; Jiang, Ren, and Lin 2023). Our approach differs from these prior works as the various LLM agents we utilize become experts in the environments from an embodied exploration phase and we show it can be used with off-the-shelf exploration methods. We believe that our approach can be considered a variation of Mixture-of-Experts (MoE) (Eigen, Ranzato, and Sutskever 2013), as early MoE methods relied on training ML models like SVM to learn the weightings of different experts, as our CAM framework does. Our main difference is that the MoE module also takes into account answers from embodied agents explore the environment.
Figure 2: Overview of experimentation pipeline. [LEFT] We have an LLM answer a collection of embodied questions about the Matterport3D environment using the observations collected and stored by a specific agent during an exploration phase. The answers from the LLM are stored offline. [RIGHT] In the central network training phase, we select multiple sets of answers from different agents to train the classifier with the ground-truth answer with supervised learning.
Multi-LLM System with Central Answer Model (CAM) for EQA
Problem Formulation
In this section, we provide a mathematical formulation of our specific EQA setting to characterize the EQA problem we are tackling in this work. We consider a tuple of $(Q,S^{K})$, where $Q$ represents the set of queries, where each sample is a vector $\left[o,r,y\right]$ with object $o$ and room $r$ representing the question “Is there an $o$ in the $r$?”, and $y\in\{0,1\}$ is the binary ground-truth answer to the question. $S$ is the set of $K$ embodied agents that independently answer the questions. Let $s_{k}^{(o,r)}\in\{0,1\}$ be the independent answer to query $(o,r)$ of the $k$-th agent. Let $f$ be a function such that $f\left(o,r,\cup_{k=1}^{K}s_{k}^{(o,r)}\right)=\hat{y}\in\{0,1\}$. Given the query context $(o,r)$ and the independent agent answers to the context, $f$ should produce a final binary prediction to whether or not object $o$ is in room $r$. In this work, we present CAM as a viable method of the function $f$ over baselines majority voting and debating. We now describe CAM in the next section.
Our Approach
We implement a two-step procedure that comprises a question-answering phase, and a central model training phase. The question-answering phase is an agent-specific step while the classification-training step relies on combining the intermediate results of the question-answering phases from various agents. Before starting the question-answering phase, the agents explored the environment to gather observations. Later, we examine the performance of our approach when using a LLM-based SOTA method for exploration. We describe both phases below.
Question-Answering Phase
In this phase, we generate a dataset of binary embodied questions specific to a Matterport 3D environment. Matterport3D is a large-scale dataset popular for scene-understanding tasks and embodied AI research, containing RGB-D panoramic views of various building environments like homes. The data is stored as a graph for each environment where the nodes are different physical locations within the home environment that an embodied agent can occupy. We use a topological graph environment because we are only concerned with high-level exploration policy in the house environment. Our approach can be used with other similar environments such as HM3D (Ramakrishnan et al. 2021).
Each LLM-based agent’s answers are based on independent observations from its exploration of the environment. To emphasize, the agents differ in their collected observations, leading to differing answers of the same query. To generate these questions we follow this template:
Do you think there is a [ITEM] in the [ROOM]? Use both common-sense reasoning about the object and room and the observation list given. Even if the relevant information is not in your observations, the answer can still be YES. Respond with YES or NO.
We then prompt the LLM to answer each question based on the stored observations of the agent. We pass the observations to the system prompt formatted as a dictionary, {ROOM NAME: [LIST OF ITEMS IN THE ROOM]}. The user prompt is a question following the template above and the system prompt is as follows: You are an embodied agent that has explored a house. I prefer definite answers to questions. The observations are: [OBSERVATION DICTIONARY]
Central Model Training Phase
To give the final answer to each question, we train a model $f$ to give a final, binary “Yes/No” output. The model input dimension is $2+K$, where $K$ is the number of agents. The first two inputs account for the vectorized question in the form of “[object, room]”, $(o,r)$. The rest of the $K$ inputs come from entering the independent answers from each of the agents chosen for the multi-LLM system, $\cup_{k=1}^{K}s_{k}^{(o,r)}$. $f$ then outputs a predicted answer $\hat{y}$ and is then optimized with the ground truth answer $y$. During inference, we input unseen questions and agent answers into the network to output a test-time answer. Due to the modularity of our approach, for each set of independent responses, we train multiple machine learning models to produce multiple central answer models to compare against baseline aggregation approaches.
Note: The classifier is overfitted to a given environment because the Multi-LLM embodied system is only for a specific environment. With our EQA variation, the agent must answer questions on what it has explored in the given environment. This overfitting allows for the personalization of our systems to the household environment. In our potential use case of a user asking a question of whether or not an object is located in a certain area of their environment, the agents should be trained and asked in the same environment.
Experimentation
We now describe our experimentation in detail. With these experiments, we aim to 1) measure the accuracy of CAM for EQA and show its performance over other aggregation baselines (Results section); 2) and 3) provide a feature importance analysis to highlight the benefit of utilizing independent agent responses in our approach (Feature Importance Analysis Section) In all experiments we use GPT-4-turbo.
Query Generation
We extract the embodied question bank from the Matterport3D environments. For each scan, we extract an offline list of observed items detected in each image of the environment using the GLIP object detection algorithm (Li et al. 2022). The items space observed by GLIP can be found in the Appendix in Table 6.
For each item and room combination, we formulate the question using the template described in Section Question-Answering Phase. Because these questions were extracted based on the observed items in the scan, we know the ground truth answer is “Yes”.
To create questions with ground-truth answers “No”, for each query generated with a “Yes” ground-truth, we formulate a question with the same item but with a random room that is not within the subset of rooms that contains that specific item.
Note: Our setup is highly dependent on the object detector used. Error can thus be introduced by the object detection with GLIP and then the LLM answer generation. Some outputs of GLIP included toilets in the living room and a TV in the hallway. When the LLM agent has a (room, item) in its observation memory in the system prompt, it will usually say Yes if asked about said room and item, but when it does not have it in memory, the questions are harder to answer, leading to error in answers generated by the independent LLM agent. More error analysis on agent answers will be studied in further in the Feature Importance Analysis section.
Evaluation
We use a SOTA algorithm, LGX (Dorbala, Mullen Jr, and Manocha 2023) for the exploration phase.
Metric. To measure and compare the performances of our CAM approach against baseline aggregation approaches, we calculate the inference time accuracy, or correctness, of the model:
$\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}_{y_{i}=\hat{y}_{i}},$ (1)
where $y_{i}$ is the ground truth label of the test question, and $\hat{y}_{i}$ is the final answer prediction from CAM. We produce CAM models with various machine learning algorithms: Neural Network (NN), Random Forest (RF), Decision Tree (DT), XGBoost (XG), RBF SVM (SVM), Linear SVM (SVM-L), Logistic Regression (LR).
The NN was implemented with Pytorch using 3 linear layers. The dimension were 16, 8, 1 in that order with a Sigmoid activation function at the end to ensure the output was in between 0 and 1. XGBoost used the default settings from the xgboost library package. RF from the sklearn library used $1000$ trees, and all other methods used the default settings from sklearn.
Baselines
We compare these different variations of CAM against two other forms of the aggregation methods used for ensemble LLM methods. Neither method is learning-based and requires no training set. We describe them below:
  • Majority Vote (MV): For each test question, we take the majority vote of the independent answers from the agents of the Multi-LLM system. We do not directly take into account room or object with a majority vote
  • Debate: For each test question, we have the agents of the Multi-LLM system participate in a turn-based discussion where they relay their thoughts to one another for 2 rounds. To initial the debate we provide each LLM-based agent context with the following system prompt.
    System prompt: You are an embodied agent in a house. Your id is [INDEX]. Here are your observations from exploring the house: [OBSERVATION DICTIONARY]. Your initial answer to whether or not there was a [OBJECT] in the [ROOM] was [INITIAL ANSWER]. Other agents may have different answers. Please debate with the other agents to come to a consensus. Here is the conversation history: [CONVERSATION HISTORY]
    The last sentence is replaced with This is the beginning of the conversation. at the start of the debate when no conversation history is present. Below, we provide the user prompt to get the agent’s response for its turn during that round.
    User prompt: It is your turn to speak. Use your observation and conversation history to help.
    When prompted for the final answer at the end of the debate. The user prompt is Please give your final ”Yes/No” answer. Give a definite answer.
We then collect the final answers from each agent and take the majority vote as the final prediction.
Language-Guided Exploration
Figure 3: LGX + CAM; Performance of CAM with $3$ independent agent that performed LGX for collecting observations. We use a 90%-10% train-test random split. We experiment on a Matterport environment with 215 nodes. We report the mean and standard deviation over $5$ trials. Against both baselines, all CAM methods outperform during test-time inference, with XGBoost having the highest accuracy.
Accuracy (%)
CAM-based Methods (Ours) Baselines
Environment Size (num of nodes) NN RF DT XG SVM SVM-L LR MV Debate
20 65.56 63.33 73.33 73.33 64.44 43.33 53.33 58.89 58.89
43 61.00 63.00 86.00 82.00 62.00 60.00 64.00 41.00 40.00
53 50.83 68.33 78.75 81.25 64.17 50.83 50.83 44.58 44.17
58 50.00 74.00 84.00 87.00 74.00 55.00 53.00 54.00 53.00
78 65.00 70.45 85.45 86.36 70.45 57.27 55.91 42.27 41.82
114 64.06 77.19 81.25 87.50 70.94 55.63 60.31 51.88 47.19
145 65.33 75.33 83.00 87.67 69.00 58.33 59.33 46.67 45.33
215 66.33 72.67 79.67 88.33 69.67 56.00 56.33 42.00 43.00
Accuracy (%) with Malicious Agent
CAM-based Methods (Ours) Baselines
Environment Size (num of nodes) NN RF DT XG SVM SVM-L LR MV Debate
20 61.11 62.22 73.33 73.33 64.44 44.44 53.33 61.11 58.89
53 60.42 68.33 78.75 81.25 64.17 49.58 50.83 46.67 44.17
145 68.00 75.33 83.00 87.67 69.00 59.33 59.33 47.67 45.67
215 64.33 73.67 79.67 88.33 69.67 56.33 56.33 43.67 43.00
Table 2: Performance measured in accuracy (Eq 1) of CAM with $3$ independent agent that performed LGX for collecting observations. We use a 90%-10% train-test random split. We experiment in $8$ different environments over $5$ trials. Against both baselines, all CAM methods outperform during test-time inference, with XGBoost having the highest accuracy. We repeated the experiment in $4$ of the environments but turned one of the agent malicious by providing responses opposite to what it believes are the correct answers. In cases where one agent is malicious, our CAM method is still better suited than MV and debate.
We have all the LLM-based agents explore the household environment for a set amount of time before answering questions. In the exploration phase, each subordinate agent is set in the Matterport house scan graph and performs an augmented variation of Language-Guided Exploration (LGX) as proposed by Dorbala, Mullen Jr, and Manocha (2023). In traditional LGX, the LLM agent observes the immediate household items around itself and chooses which one to navigate to find a specific target item, forming a chain of “hops” to the item.
In our variation, we instead provide the LLM with a system prompt that encourages the agent to explore as much of the house as possible. You are an embodied agent in a house. You want to aggressively explore the house and you want to find as many unique objects as you can.
The agent then travels from node to node in the house graph. For each timestep, we document the timestep, node, and items observed. When the agent reaches a node we can then access the stored results and add them to the list of observed items. We have the $3$ agents perform $10$ steps of this augmented LGX and store the observations. The rest of the experiment process follows the same setup as described in the Approach Section with the Question-Answering and Central Model Training phases. We use a $90\%$-$10\%$ random train-test split over $5$ seeds (0 to 4 are used as the seeds). Experiments are ran on an Apple M1 Pro and macOS 14.5.
Results: The performance of the various CAM methods and baselines in the $8$ graph environments is shown in Table 2. We test in environments of different sizes (measured in number of nodes) to showcase the generalization of the CAM approach across different size complexities. From the table, we see CAM model consistently outperforms both baselines in accuracy (Eq 1. We report the average accuracy over $5$ seeds. We see that the non-linear methods consistently achieve higher accuracy than the linear methods, suggesting that the dataset contains no hyperplane separating the queries with “Yes” and “No” ground-truth answers. Specifically, XGBoost attains the highest accuracy in all but one environment, suggesting that the boosted ensemble method can best fit the generated query-answer data.
Interestingly, in all but one of the environments, debating does slightly worse on average than MV. This degradation occurs due to agents who were initially correct being swayed by incorrect agents during conversation. Overall, our Multi-LLM system can be utilized in conjunction with an LLM-based exploration method, emphasizing its practicality in real-world use cases.
Method Agent 1 Agent 2 Agent 3
XG 38 $\pm$ 8.3 38 $\pm$ 7.0 38 $\pm$ 6.2
MV 97 $\pm$ 0.1 95 $\pm$ 0.1 96 $\pm$ 1.6
debate 97 $\pm$ 1.4 93 $\pm$ 3.1 91 $\pm$ 3.6
Table 3: Agreement (%) between each LGX agent and the final answers of 1) XGBoost-based CAM (XG), 2) Majority Vote (MV) and 3) debate in the $215$-node environment. We report the average and standard deviation over the $5$ trials. We see that XGBoost learns not to agree with any of the agents to achieve its high accuracy, suggesting more exploration or better object detection is needed.
Agreement: We analyze the agreement between each agent and the XGBoost model (the best performing CAM-based method), majority vote, and debate. We measure agreement between an agent and each central module method by the percentage of similar answers:
$Agr(s_{k},f)=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}_{s_{k}^{i}=\hat{y}_{i}},$ (2)
where $i$ is index of the query parameterized by $(o,r)$. Table 3 reports the agreement over $5$ trials and we see that XGBoost has learned not to agree with the agents as much as MV and debate to attain its high accuracy. This low agreement could stem from the errors introduced by the object detection of GLIP and the question-answering of each individual agent. We further analyze the independent answers impact on training in the Feature Importance Analysis Section.
Query Subject $\downarrow$ Object Kitchen Dining Room Bedroom Office
Kitchen 27.50 22.50 27.50 27.50 22.50
Dining Room 17.50 12.50 12.50 17.50 12.50
Bedroom 25.00 25.00 25.00 25.00 25.00
Office 12.50 5.00 5.00 12.50 7.50
Table 4: This table displays the feature importance of different decision tree-based CAM models. We see in cases where the queries only pertaining to the dining room (third row) that responses based only on dining room observation (fourth column) is not the most important feature. This insight suggests that CAM learned not to rely on that LLM agent.
Malicious Agent: We also show the benefits of CAM where one of the three agents is malicious. To turn the agent malicious we invert its independent answers so that all “Yes” answers turn to “No” and vice versa. We then train all CAM methods and execute baselines regularly. Thus, the malicious agent contributes an answer it believes to be incorrect. From Table 2, we see that our CAM methods still outperform the MV and debate baselines. We see that the debate baseline again does worse the majority vote, showing that the incorrect agent was able to sway the overall decision that was initially going to be correct. In the next section, we provide a feature importance of CAM.
Feature Importance Analysis
In the context of machine learning, feature importance is a metric by which input impacts the model’s final answer. For our Multi-LLM system, feature importance can be used to examine if any agent response or query context (i.e. room or object) is relied heavily upon for training the central answer model. To examine which features hold the most weight in the final decision of CAM model, we look into studying the variable, or feature importance (Breiman 2001; Fisher, Rudin, and Dominici 2018), of the CAM model. We measure feature importance using permutation feature importance (PFI). We define the PFI the $i$-th feature as,
$PFI(i)=|a_{\text{val}}(m)-a_{\text{val}}^{i}(m)|,$ (3)
where $a_{\text{val}}(m)$ is the model $m$ accuracy measured by Equation 1 of the original validation set, and $a_{\text{val}}^{i}(m)$ is the accuracy with a modified validation set where the $i$-th feature column is randomly permuted. Intuitively, the more important a variable, the more the accuracy should change if the values are randomly shuffled because the variable is highly correlated with the target variable. In our Multi-LLM system for EQA, an agent that gives nearly-deterministic answers (e.g. “No” for almost all answers) would have a PFI close to 00.
To examine feature importance, we train multiple DT-based CAM models, each with a different filtered query dataset from the $215$-node environment. To better highlight and focus on the feature importance aspect, we have an oracle exploration where we explicitly say which rooms and items an agent has in its observations memory when answering the questions. This feature importance analysis could be applied to dynamic agents that move around with LGX.
In each training, the features comprise the room and object name to represent the query and $4$ independent responses from LLMs that had observations either from the kitchen, dining room, bedroom, or office. The dataset however will only consist of rows with a given “Room” value. The model is trained and tested, and we then calculate feature importance using Equation 3.
Table 4 summarizes our feature importance findings. The leftmost column specifies the room the dataset was filtered for, and the other columns show the average PFI of the object feature and independent responses for that filtered dataset over 5 trials. We see that in many cases, like when trained only on the dining room queries, the corresponding feature (i.e. response based only on dining room observation) does not have the highest PFI. We analyzed the offline LLM-generated answers based only on dining room observations and out of the $600$ responses, $549$ are “No”, or $91.5\%$. Thus, the CAM model has learned not to rely on these answers as much as the other responses even though they are based only on observations from the room in question. This outcome demonstrates how analyzing feature importance can help examine individual members of a multi-agent setup for error analysis and further supports our approach utilizing independent responses. Please see the Appendix for a visualization of the Decision Tree to highlight the benefits of the interpretable ML-based CAM methods.
Conclusion and Limitations
In this paper, we tackle a variation of Embodied Question Answering with a Multi-LLM framework that aggregates the independent responses of LLM-based agents and produces a final answer. While other methods have used a central module to aggregate independent agent answers for QA tasks, we specifically investigate this method in the context of embodied exploration in a household environment. We propose a central answer model (CAM) that learns to weigh the responses of each agent to decide on a single answer to the given query. We show that various machine learning methods are suitable to train CAM and consistently outperform non-learning-based aggregation methods for ensemble outputs, specifically majority voting and multi-LLM debating. A key benefit of our system is that it requires no communication during inference, reducing time costs. Furthermore, our approach can be combined with SOTA methods for exploration in unknown households. We also show that CAM is more robust to malicious agents than the majority vote and debate baslines. We end this work by providing a feature importance analysis to study the effect of individual agents and target object specification on the final answer. We hope that this work will encourage research in embodied Multi-LLM for QA as LLM-based agents become more household.
This work has several limitations. First, the ground truth labels to train CAM can be difficult to obtain, especially if the household environment is constantly changing with non-stationary items. We leave it as future work to adapt this Multi-LLM setup for EQA in dynamic household environments. Second, our queries are all objective binary “Yes/No” questions. To better enhance practicality, future work should look into aggregation methods for situational, subjective, and non-binary questions.
Appendix A Appendix
Appendix B Code and Dataset
We provide an anonymous link to our code  [^1] https://anonymous.4open.science/r/CAM-embodied-llm-54CE/. The query dataset extracted is also within the submitted supplementary materials folder.
Appendix C Multi-LLM vs LLM
To motivate multiple LLM agents we provide a comparison between 1 agent and 3 agents in Table 5.
Environment Size Accuracy (%) of 1 Agent Accuracy (%) of 3 Agents (MV)
$53$ nodes $41.67$ $42.00$
$215$ node $44.17$ $44.58$
Table 5: Rise of accuracy with multiple dynamic agents
Appendix D Additional Visualization of Results
In this section, to focus on the aggregation abilities of CAM, we have an oracle exploration where we explicitly say which rooms and items an agent has in its observations memory when answering the questions.
Below we provide additional visualizations of the results of our learning-based CAM methods against MV and debating on other Matterport environments. Like Table 2, we use independent answers of $3$ LLMs explored with LGX. Majority of the CAM methods consistently outperform the baseline majority vote (MV) and debate methods, with Decision Tree (DT) and XGBoost (XG) being the highest-performing methods.
Figure 4: 20 nodes
Figure 5: 43 nodes
Figure 6: 53 nodes
Figure 7: 58 nodes
Figure 8: 78 nodes
Figure 9: 114 nodes
Figure 10: 145 nodes
Appendix E Decision Tree Visualization
In this section, we showcase the potential interpretability of our learning-based CAM framework by visualizing the decision tree trained in Section Feature Importance Analysis. This tree in Figure 11 was trained on queries only pertaining to the kitchen and was trained on independent answers from four different LLM agents. One agent had only observations from kitchen; another only dining room; another bedroom; and the last one only had office observations. We see that the main feature used was actually the object query. We provide the object encoding labels in Table 6.
Figure 11: Visualization of Decision Tree from Section Feature Importance Analysis.
Object Label Encoding $o$
appliance 0
armchair 1
bathtub 2
bed 3
board 4
bookcase 5
books 6
cabinet 7
chair 8
clothes 9
counter 10
curtain 11
cushion 12
desk 13
dresser 14
dumbells 15
hairbrush 16
headphones 17
jumprope 18
mirror 19
mug 20
nightstand 21
phone 22
picture 23
plant 24
playing cards 25
refrigerator 26
sink 27
sofa 28
table 29
television 30
toilet 31
toothbrush 32
towel 33
tv remote 34
wallet 35
water bottle 36
whiteboard 37
window 38
wristwatch 39
Table 6: Mapping of object labels to encodings $o$