### v5

# Multi-LLM QA with Embodied Exploration

Abstract
Large language models (LLMs) have grown in popularity due to their natural language interface and pre trained knowledge, leading to rapidly increasing success in question-answering (QA) tasks. More recently, multi-agent systems with LLM-based agents (Multi-LLM) have been utilized increasingly more for QA. In these scenarios, the models may each answer the question and reach a consensus or each model is specialized to answer different domain questions. However, most prior work dealing with Multi-LLM QA has focused on scenarios where the models are asked in a zero-shot manner or are given information sources to extract the answer. For question answering of an unknown environment, embodied exploration of the environment is first needed to answer the question. This skill is necessary for personalizing embodied AI to environments such as households. There is a lack of insight into whether a Multi-LLM system can handle question-answering based on observations from embodied exploration. In this work, we address this gap by investigating the use of Multi-Embodied LLM Explorers (MELE) for QA in an unknown environment. Multiple LLM-based agents independently explore and then answer queries about a household environment. We analyze different aggregation methods to generate a single, final answer for each query: debating, majority voting, and training a central answer module (CAM). Using CAM, we observe a $46\%$ higher accuracy compared against the other non-learning-based aggregation methods. We provide code and the query dataset for further research.
Multi-LLM QA with Embodied Exploration
Bhrij Patel University of Maryland, College Park                        Vishnu Sashank Dorbala University of Maryland, College Park Dinesh Manocha University of Maryland, College Park                        Amrit Singh Bedi University of Central Florida
1 Introduction
Large Language Models (LLMs) (Achiam et al., 2023) become more embedded into society, with one of the most popular tasks being question-answering (QA) (Rawal et al., 2024; Kamalloo et al., 2023). Further research has investigated utilizing an ensemble of LLM models (Multi-LLM) for the ability to discuss (Li et al., 2023; Chen et al., 2023b), answer questions from large data sources and route questions to models with specialized knowledge (Mixture of Experts) (Li et al., 2024).
However, in these prior works on ensemble LLMs for QA, the questions are asked in a zero-shot manner or information is provided to the models. For QA of a physical environment, an embodied agent must first explore an environment, such as a household, to answer questions based on its observations. It has gained considerable interest in the past few years due to its applications for in-home robots and personalized assistants Das et al. (2018); Gordon et al. (2018); Duan et al. (2022a); Sima et al. (2023); Luo et al. (2023); Dorbala et al. (2024a); Zhu et al. (2023). Possessing zero-shot common-sense reasoning of relational understandings of household items is crucial for LLMs to perform embodied exploration in unknown environments Dorbala et al. (2023, 2024b); Majumdar et al. (2024).
Figure 1: Overview of our MELE framework. LLM-based agents that have explored the dynamic household are independently asked a binary “Yes/No” question and must answer based on their observations. Only one of the agents has the correct answer (“No”) while the other two agents are wrong. The aggregator function then takes in these responses to decide a final response.
In existing works, only a single agent explores the environment with a limited field of view, inducing low coverage and high exploration costs. In contrast, for an LLM-based embodied agent to succeed in EQA, efficient exploration of the environment is needed as the questions will center around the state of the environment Das et al. (2018); Yu et al. (2019); Zhu et al. (2023); Majumdar et al. (2024). The rise of Embodied Artificial Intelligence (EAI) Duan et al. (2022b) brings up the possibility of co-habited spaces with multiple agents and humans working in conjunction Puig et al. (2023).
With this trend, we believe investigating the abilities of multiple embodied LLM agents to answer questions on their environment becomes increasingly crucial for the personalization of AI. We consider a similar co-inhabited setup and propose using a team comprising multiple embodied LLM-based explorers. An ensemble system of explorers allows for simultaneous exploration of the environment in the same amount of time.
However, there are also challenges when introducing a multi-agent setup. A key challenge in a multi-agent setup is with conflicting responses when agents are asked the same query. Since LLMs are blackboxes capable of generating free-form natural language, their responses to the same query may significantly vary Gal and Grosz (2022); Wang et al. (2024b). In an embodied setting, such responses are further skewed by partial observations due to limited exploration. Prior work in ensemble LLM methods tackle this issue using weighted sum or voting schemes Li et al. (2024); Wei et al. (2022); Wang et al. (2022); Yang et al. (2024) and consensus-reaching debates between the LLM agents Chen et al. (2023a); Du et al. (2023); Xiong et al. (2023); Guo et al. (2024a); Chan et al. (2023); Li et al. (2023). However, such methods can be vulnerable to scenarios where incorrect answers from malicious or poor-performing agents can influence the overall decision Amayuelas et al. (2024).
Main Results: To investigate the use of ensemble LLM for QA with embodied exploration, we apply a system of Multi-Embodied LLM Explorers (MELE) to solve this embodied task. We instruct the agents to answer based on their observations a set of binary embodied questions about the household. We then use various aggregation methods to output a final answer based on the independent responses from the agents: majority voting, debating, and training a supervised central answer model (CAM). Majority voting and debating are common aggregation methods in existing literature and we propose investigating training CAM for more robustness to adversarial or incorrect agents. Our novel contributions include:
  • Multi-LLM QA Framework with Embodied Navigation: We use multiple independent embodied LLM agents that explore the environment individually generate a response and use various aggregator functions to predict a final response.  [^1] Link to anonymous repository: https://anonymous.4open.science/r/CAM-embodied-llm-54CE/ We depart from prior Embodied Question-Answering (EQA) Das et al. (2018) setups where the agents explore after being asked a question. Rather we have the agents explore before being asked a set of questions on their environment. We generate queries for various Matterport 3D Chang et al. (2017) environments. We train CAM on the labeled query dataset with multiple machine learning methods and show that the best CAM-based method can achieve up to 46% higher accuracy than a majority vote and debate aggregation methods. While prior work has used a central answer module to output a final answer from a mixture of answers from multiple LLM experts Puerto et al. (2023); Si et al. (2023); Jiang et al. (2023), we are the first to apply the framework to embodied agents that must first explore an environment.
  • Integration with Exploration Systems: We evaluate our multi-agent framework on data gathered using an LLM-based exploration method Dorbala et al. (2023) on multiple agents in Matterport3D, demonstrating our system can work in conjunction with SOTA exploration methods in unknown environments.
  • Examining Agent and Feature Reliance: We present a feature importance analysis of CAM reliance on each independent agent using the metric, permutation feature importance (PFI) Breiman (2001); Fisher et al. (2018). We also look at the agreement between the individual agent and the final answer of the central answer model and the baselines. Similarly to Si et al. (2023), some of our methods for the central answer model are interpretable and provide a Decision Tree visualization of one of our models.
2 Related Works
2.1 Embodied Question Answering and Exploration
Embodied Question Answering (EQA) was first introduced by Das et al. (2018) and has become a popular task for robotics (Gordon et al., 2018, 2019; Duan et al., 2022a; Sima et al., 2023; Wijmans et al., 2019; Chen et al., 2019; Luo et al., 2019), where traditionally an agent is given a query such as “was there a plant in the living room?’ and explores its environment, in this case a house, to answer the question. Zhu et al. (2023) had alternated between an exploration phase and then a question phase and repeated the process until the query was answered. Further work has extended EQA to scenarios where the final answer revolves around multiple objects Yu et al. (2019), and where a target object is not explicitly stated in the question Tan et al. (2023). Tan et al. (2020) introduced multi-agent exploration and cooperation for question-answering in interactive environments. Dorbala et al. (2024a) provided a dataset of subjective or situational queries and was the first to use a generative approach for query creation, and Majumdar et al. (2024) designed the first open-vocabulary EQA dataset to test and benchmark foundation models on this task. Recently, Singh et al. (2024) studied the zero-shot LLM performance on visual QA with indoor house scenes. However, the LLM is given the scenes without any exploration of the environment. We evaluate our Multi-LLM approach with a SOTA LLM-based exploration method for observation gathering introduced by Dorbala et al. (2023).
Note: In our work, we deviate from the traditional EQA approach. We first have an exploration gathering phase, then answer a set of questions in the same environment. We do so for two reasons: 1) to provide a setup that better resembles prior work in LLM-based MoE QA (Li et al., 2024) 2) we believe this setup is better suited for the scenario where the user asks a question about the items in their homes and quickly needs an answer.
2.2 Systems of Multiple LLM-based Agents
Research in LLM-based systems have grown in popularity with the rise of foundation models for areas such as task planning Chen et al. (2023c), application development Wu et al. (2023), and chatbots Duan et al. (2023). Extensive studies have focused on their abilities to cooperate. Prior work has used debate between LLM models to reach a consensus (Chan et al., 2023; Du et al., 2023; Guo et al., 2024a; Chen et al., 2023b). Chen et al. (2023a) analyzes the consensus-seeking debate between multiple LLM agents with different personalities, and Amayuelas et al. (2024) investigates the susceptibility of a LLM-based discussion to an adversarial agent that purposefully tries to lead an incorrect answer choice. For embodied AI, Guo et al. (2024b) have shown that multiple LLM-based agents can cooperate in teams by designating a leader for increased task efficiency. Other work have looked into using voting metrics of independent LLM-agents to arrive at a final answer (Li et al., 2024; Wei et al., 2022; Wang et al., 2022). Our work aims to remove the need for communication by training a central classifier on independent answers from multiple agents. Training a classifier in a supervised manner helps our approach become robust to unreliable agents that may influence overall answers in voting schemes.
A central architecture to output a final answer from multiple answers from different LLM experts has been explored previously Puerto et al. (2023); Si et al. (2023); Jiang et al. (2023). Our approach differs from these prior works as the various LLM agents we utilize become experts in the environments from an embodied exploration phase and we show it can be used with off-the-shelf exploration methods. Furthermore, Wang et al. (2024a) introduced the idea of a Mixture of Agents to enhance task performance of LLMs. We specifically look at a mixture of embodied agents that explore a unknown environment.
3 Multi-Embodied LLM Explorers
3.1 Problem Formulation
In this section, we provide a mathematical formulation of our specific EQA setting to characterize the EQA problem we are tackling in this work. We consider a tuple of $(Q,S^{K})$, where $Q$ represents the set of queries, where each sample is a vector $\left[o,r,y\right]$ with object $o$ and room $r$ representing the question “Is there an $o$ in the $r$?”, and $y\in\{0,1\}$ is the binary ground-truth answer to the question. $S$ is the set of $K$ embodied agents that independently answer the questions. Let $s_{k}^{(o,r)}\in\{0,1\}$ be the independent answer to query $(o,r)$ of the $k$-th agent. Let $f$ be a aggregation function such that $f\left(o,r,\cup_{k=1}^{K}s_{k}^{(o,r)}\right)=\hat{y}\in\{0,1\}$. Given the query context $(o,r)$ and the independent agent answers to the context, $f$ should produce a final binary prediction of whether object $o$ is in room $r$.
3.2 Our Approach
We implement MELE as a two-step procedure that comprises a question-answering phase, and a aggregation phase. The question-answering phase is an agent-specific step while the classification-training step relies on combining the intermediate results of the question-answering phases from various agents. Before starting the question-answering phase, the agents explored the environment to gather observations. Later, we examine the performance of MELE when using a LLM-based SOTA method for exploration.
3.3 Question-Answering Phase
In this phase, we generate a dataset of binary embodied questions specific to a Matterport 3D environment. Matterport3D is a large-scale dataset popular for scene-understanding tasks and embodied AI research, containing RGB-D panoramic views of various building environments like homes. The data is stored as a graph for each environment where the nodes are different physical locations within the home environment that an embodied agent can occupy. We use a topological graph environment because we are only concerned with high-level exploration policy in the house environment. Our approach can be used with other similar environments such as HM3D Ramakrishnan et al. (2021).
Each LLM-based agent’s answers are based on independent observations from its exploration of the environment. To emphasize, the agents differ in their collected observations, leading to differing answers of the same query. To generate these questions we follow this template:
Do you think there is a [ITEM] in the [ROOM]? Use both common-sense reasoning about the object and room and the observation list given. Even if the relevant information is not in your observations, the answer can still be YES. Respond with YES or NO.
We then prompt the LLM to answer each question based on the stored observations of the agent. We pass the observations to the system prompt formatted as a dictionary, {ROOM NAME: [LIST OF ITEMS IN THE ROOM]}. The user prompt is a question following the template above and the system prompt is as follows: You are an embodied agent that has explored a house. I prefer definite answers to questions. The observations are: [OBSERVATION DICTIONARY]
3.4 Aggregation Phase
We now describe the different methods used for aggregation function $f$.
3.4.1 Non-learning-based Aggregation
We first describe the non-learning based methods used commonly in ensemble LLM literature.
  • Majority Vote (MV): For each test question, we take the majority vote of the independent answers from the agents of the Multi-LLM system. We do not directly take into account room or object with a majority vote
  • Debate: For each test question, we have the agents of the Multi-LLM system participate in a turn-based discussion where they relay their thoughts to one another for 2 rounds. To initialize the debate we provide each LLM-based agent context with the following system prompt.
    System prompt: You are an embodied agent in a house. Your id is [INDEX]. Here are your observations from exploring the house: [OBSERVATION DICTIONARY]. Your initial answer to whether or not there was a [OBJECT] in the [ROOM] was [INITIAL ANSWER]. Other agents may have different answers. Please debate with the other agents to come to a consensus. Here is the conversation history: [CONVERSATION HISTORY]
    The last sentence is replaced with This is the beginning of the conversation. at the start of the debate when no conversation history is present. Below, we provide the user prompt to get the agent’s response for its turn during that round.
    User prompt: It is your turn to speak. Use your observation and conversation history to help.
    When prompted for the final answer at the end of the debate. The user prompt is Please give your final "Yes/No" answer. Give a definite answer. We then collect the final answers from each agent and take the majority vote as the final prediction.
3.4.2 Central Answer Model (CAM)
To give the final answer to each question, we train a model $f$ to give a final, binary “Yes/No" output. The model input dimension is $2+K$, where $K$ is the number of agents. The first two inputs account for the vectorized question in the form of “[object, room]”, $(o,r)$. The rest of the $K$ inputs come from entering the independent answers from each of the agents chosen for the multi-LLM system, $\cup_{k=1}^{K}s_{k}^{(o,r)}$. $f$ then outputs a predicted answer $\hat{y}$ and is then optimized with the ground truth answer $y$. During inference, we input unseen questions and agent answers into the network to output a test-time answer. Due to the modularity of our approach, for each set of independent responses, we train multiple machine learning models to produce multiple central answer models to compare against baseline aggregation approaches.
Note: The classifier is overfitted to a given environment because the Multi-LLM embodied system is only for a specific environment. With our EQA variation, the agent must answer questions on what it has explored in the given environment. This overfitting allows for the personalization of our systems to the household environment. In our potential use case of a user asking a question of whether or not an object is located in a certain area of their environment, the agents should be trained and asked in the same environment.
We produce CAM models with various machine-learning algorithms: Neural Network (NN), Random Forest (RF), Decision Tree (DT), XGBoost (XG), RBF SVM (SVM), Linear SVM (SVM-L), and Logistic Regression (LR). The NN was implemented with Pytorch using 3 linear layers. The dimension were 16, 8, 1 in that order with a Sigmoid activation function at the end to ensure the output was in between 0 and 1. XGBoost used the default settings from the xgboost library package. RF from the sklearn library used $1000$ trees, and all other methods used the default sklearn settings.
4 Experimentation
We now describe our experimentation in detail. With these experiments, we aim to 1) measure the accuracy of CAM for EQA and show its performance over other aggregation baselines (Results section); 2) and 3) provide a feature importance analysis to highlight the benefit of utilizing independent agent responses in our approach (Feature Importance Analysis Section). In all experiments we use GPT-4-turbo.
4.1 Query Generation
We extract the embodied question bank from the Matterport3D environments. For each scan, we extract an offline list of observed items detected in each image of the environment using the GLIP object detection algorithm Li et al. (2022). The items space observed by GLIP can be found in the Appendix in Table 5.
For each item and room combination, we formulate the question using the template described in Section 3.3. Because these questions were extracted based on the observed items in the scan, we know the ground truth answer is “Yes".
To create questions with ground-truth answers “No", for each query generated with a “Yes" ground-truth, we formulate a question with the same item but with a random room that is not within the subset of rooms that contains that specific item.
Note: Our setup is highly dependent on the object detector used. Error can thus be introduced by the object detection with GLIP and then the LLM answer generation. Some outputs of GLIP included toilets in the living room and a TV in the hallway. When the LLM agent has a (room, item) in its observation memory in the system prompt, it will usually say Yes if asked about said room and item, but when it does not have it in memory, the questions are harder to answer, leading to error in answers generated by the independent LLM agent. More error analysis on agent answers will be studied further in the Section 5.
4.2 Evaluation
We use a SOTA algorithm, LGX Dorbala et al. (2023) for the exploration phase.
Metric. To measure and compare the performances of our various MELE aggregation methods, we calculate the inference time accuracy, or correctness, of the model:
$\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}_{y_{i}=\hat{y}_{i}},$ (1)
where $y_{i}$ is the ground truth label of the test question, and $\hat{y}_{i}$ is the final answer prediction from the given aggregation method.
4.3 Language-Guided Exploration
Accuracy (%)
CAM-based Methods (Ours) Baselines
Environment Size (num of nodes) NN RF DT XG SVM SVM-L LR MV Debate
20 65.56 63.33 73.33 73.33 64.44 43.33 53.33 58.89 58.89
43 61.00 63.00 86.00 82.00 62.00 60.00 64.00 41.00 40.00
53 50.83 68.33 78.75 81.25 64.17 50.83 50.83 44.58 44.17
58 50.00 74.00 84.00 87.00 74.00 55.00 53.00 54.00 53.00
78 65.00 70.45 85.45 86.36 70.45 57.27 55.91 42.27 41.82
114 64.06 77.19 81.25 87.50 70.94 55.63 60.31 51.88 47.19
145 65.33 75.33 83.00 87.67 69.00 58.33 59.33 46.67 45.33
215 66.33 72.67 79.67 88.33 69.67 56.00 56.33 42.00 43.00
Accuracy (%) with Malicious Agent
CAM-based Methods (Ours) Baselines
Environment Size (num of nodes) NN RF DT XG SVM SVM-L LR MV Debate
20 61.11 62.22 73.33 73.33 64.44 44.44 53.33 61.11 58.89
53 60.42 68.33 78.75 81.25 64.17 49.58 50.83 46.67 44.17
145 68.00 75.33 83.00 87.67 69.00 59.33 59.33 47.67 45.67
215 64.33 73.67 79.67 88.33 69.67 56.33 56.33 43.67 43.00
Table 1: Performance measured in accuracy (Eq 1) of MELE with $3$ independent agent that performed LGX for collecting observations. We use a 90%-10% train-test random split. We experiment in $8$ different environments over $5$ trials. Against both baselines, all CAM methods outperform during test-time inference, with XGBoost having the highest accuracy. We repeated the experiment in $4$ of the environments but turned one of the agent malicious by providing responses opposite to what it believes are the correct answers. In cases where one agent is malicious, our CAM method is still better suited than MV and debate.
Figure 2: Performance of MELE methods with $3$ independent agent that performed LGX for collecting observations. We use a 90%-10% train-test random split. We experiment on a Matterport environment with 215 nodes. We report the mean and standard deviation over $5$ trials. Against both baselines, all CAM methods outperform during test-time inference, with XGBoost having the highest accuracy.
We have all the LLM-based agents explore the household environment for a set amount of time before answering questions. In the exploration phase, each subordinate agent is set in the Matterport house scan graph and performs an augmented variation of Language-Guided Exploration (LGX) as proposed by Dorbala et al. (2023). In traditional LGX, the LLM agent observes the immediate household items around itself and chooses which one to navigate to find a specific target item, forming a chain of “hops” to the item.
In our variation, we instead provide the LLM with a system prompt that encourages the agent to explore as much of the house as possible. You are an embodied agent in a house. You want to aggressively explore the house and you want to find as many unique objects as you can.
The agent then travels from node to node in the house graph. For each timestep, we document the timestep, node, and items observed. When the agent reaches a node we can then access the stored results and add them to the list of observed items. We have the $3$ agents perform $10$ steps of this augmented LGX and store the observations. The rest of the experiment process follows the same setup as described in Section 3 with the Question-Answering and Aggregation phases. We use a $90\%$-$10\%$ random train-test split over $5$ seeds (0 to 4 are used as the seeds). Experiments were run on an Apple M1 Pro and macOS 14.5.
Results: The performance of the various CAM methods and non-learning aggregation methods in the $8$ graph environments is shown in Table 1. We test in environments of different sizes (measured in number of nodes) to showcase the generalization of the CAM approach across different size complexities. From the table, we see CAM model consistently outperforms both MV and debating in accuracy (Eq 1. We report the average accuracy over $5$ seeds. We see that the non-linear methods consistently achieve higher accuracy than the linear methods, suggesting that the dataset contains no hyperplane separating the queries with “Yes” and “No” ground-truth answers. Specifically, XGBoost attains the highest accuracy in all but one environment, suggesting that the boosted ensemble method best fits the generated query-answer data.
Interestingly, in all but one of the environments, debating does slightly worse on average than MV. This degradation occurs due to agents who were initially correct being swayed by incorrect agents during conversation. Overall, our Multi-LLM system can be utilized in conjunction with an LLM-based exploration method, emphasizing its practicality in real-world use cases.
Method Agent 1 Agent 2 Agent 3
XG 38 $\pm$ 8.3 38 $\pm$ 7.0 38 $\pm$ 6.2
MV 97 $\pm$ 0.1 95 $\pm$ 0.1 96 $\pm$ 1.6
debate 97 $\pm$ 1.4 93 $\pm$ 3.1 91 $\pm$ 3.6
Table 2: Agreement (%) between each LGX agent and the final answers of 1) XGBoost-based CAM (XG), 2) Majority Vote (MV) and 3) debate in the $215$-node environment. We report the average and standard deviation over the $5$ trials. We see that XGBoost learns not to agree with any of the agents to achieve its high accuracy, suggesting more exploration or better object detection is needed.
Agreement: We analyze the agreement between each agent and the XGBoost model (the best performing CAM-based method), majority vote, and debate. We measure agreement between an agent and each central module method by the percentage of similar answers:
$Agr(s_{k},f)=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}_{s_{k}^{i}=\hat{y}_{i}},$ (2)
where $i$ is index of the query parameterized by $(o,r)$. Table 2 reports the agreement over $5$ trials and we see that XGBoost has learned not to agree with the agents as much as MV and debate to attain its high accuracy. This low agreement could stem from the errors introduced by the object detection of GLIP and the question-answering of each individual agent. We further analyze the independent answers’ impact on training in the Section 5.
Query Subject $\downarrow$ Object Kitchen Dining Room Bedroom Office
Kitchen 27.50 22.50 27.50 27.50 22.50
Dining Room 17.50 12.50 12.50 17.50 12.50
Bedroom 25.00 25.00 25.00 25.00 25.00
Office 12.50 5.00 5.00 12.50 7.50
Table 3: This table displays the feature importance of different decision tree-based CAM models. We see in cases where the queries only pertaining to the dining room (third row) that responses based only on dining room observation (fourth column) is not the most important feature. This insight suggests that CAM learned not to rely on that LLM agent.
Malicious Agent: We also show the benefits of CAM where one of the three agents is malicious. To turn the agent malicious we invert its independent answers so that all “Yes” answers turn to “No” and vice versa. We then train all CAM methods and execute MV and debating regularly. Thus, the malicious agent contributes an answer it believes to be incorrect. From Table 1, we see that our CAM methods still outperform MV and debate. We see that the debate baseline again does worse the majority vote, showing that the incorrect agent was able to sway the overall decision that was initially going to be correct.
5 Feature Importance Analysis
In the context of machine learning, feature importance is a metric by which input impacts the model’s final answer. For our MELE system, feature importance can be used to examine if any agent response or query context (i.e. room or object) is relied heavily upon for training the central answer model. To examine which features hold the most weight in the final decision of the CAM model, we look into studying the variable, or feature importance Breiman (2001); Fisher et al. (2018), of the CAM model. We measure feature importance using permutation feature importance (PFI). We define the PFI the $i$-th feature as,
$PFI(i)=|a_{\text{val}}(m)-a_{\text{val}}^{i}(m)|,$ (3)
where $a_{\text{val}}(m)$ is the model $m$ accuracy measured by Equation 1 of the original validation set, and $a_{\text{val}}^{i}(m)$ is the accuracy with a modified validation set where the $i$-th feature column is randomly permuted. Intuitively, the more important a variable, the more the accuracy should change if the values are randomly shuffled because the variable is highly correlated with the target variable. In our Multi-LLM system for EQA, an agent that gives nearly-deterministic answers (e.g. “No” for almost all answers) would have a PFI close to 00.
To examine feature importance, we train multiple DT-based CAM models, each with a different filtered query dataset from the $215$-node environment. To better highlight and focus on the feature importance aspect, we have an oracle exploration where we explicitly say which rooms and items an agent has in its observations memory when answering the questions. This feature importance analysis could be applied to dynamic agents that move around with LGX.
In each training, the features comprise the room and object name to represent the query and $4$ independent responses from LLMs that had observations either from the kitchen, dining room, bedroom, or office. The dataset however will only consist of rows with a given “Room” value. The model is trained and tested, and we then calculate feature importance using Equation 3.
Table 3 summarizes our feature importance findings. The leftmost column specifies the room the dataset was filtered for, and the other columns show the average PFI of the object feature and independent responses for that filtered dataset over 5 trials. We see that in many cases, like when trained only on the dining room queries, the corresponding feature (i.e. response based only on dining room observation) does not have the highest PFI. We analyzed the offline LLM-generated answers based only on dining room observations and out of the $600$ responses, $549$ are “No”, or $91.5\%$. Thus, the CAM model has learned not to rely on these answers as much as the other responses even though they are based only on observations from the room in question. This outcome demonstrates how analyzing feature importance can help examine individual members of a multi-agent setup for error analysis and further supports our approach utilizing independent responses. Please see the Appendix for a visualization of the Decision Tree to highlight the benefits of the interpretable ML-based CAM methods.
6 Conclusion
In this paper, we use Multi-Embodied LLM Explorers (MELE) to understand the effects of embodied exploration of multiple independent LLM-based agents on QA task for unknown environment. While other methods in the field of Mixture of Experts have used multiple LLM-based models for QA tasks, we specifically investigate this paradigm with a dependence on embodied exploration in a household environment. We compare the accuracy of various aggregation methods: majority voting, debating, and CAM. We show that various machine learning methods are suitable to train CAM and consistently outperform the non-learning-based aggregation methods.
A key benefit of CAM is that it requires no communication during inference, reducing time costs. We also show that CAM is more robust to malicious agents than the majority vote and debate methods. We end this work by providing a feature importance analysis to study the effect of individual agents and target object specification on the final answer. We hope that this work will encourage research in embodied Multi-LLM for QA as LLM-based agents become more household.
7 Limitations and Future Work
This work has several limitations. First, the ground truth labels to train CAM can be difficult to obtain, especially if the household environment is constantly changing with non-stationary items. We leave it as future work to adapt this Multi-LLM setup for EQA in dynamic household environments. Second, our queries are all objective binary “Yes/No” questions. To better enhance practicality, future work should look into aggregation methods for situational, subjective, and non-binary questions. It would also be useful to remove the need to explicitly mention the target object in the query. Other areas directions for research could be then to apply our CAM aggregation method to question-answering tasks outside of Embodied AI such as long video understanding. Third, our results hinge on the quality of the observation data. In our work, we relied on the GLIP model. Future work should analyze the effect of scene understanding algorithms on our setup.
Appendix A Code and Dataset
We provide an anonymous link to our code  [^2] https://anonymous.4open.science/r/CAM-embodied-llm-54CE/. The query dataset extracted is also within the submitted supplementary materials folder.
Appendix B Multi-LLM vs LLM
To motivate multiple LLM agents we provide a comparison between 1 agent and 3 agents in Table 4.
Environment Size Accuracy (%) of 1 Agent Accuracy (%) of 3 Agents (MV)
$53$ nodes $41.67$ $42.00$
$215$ node $44.17$ $44.58$
Table 4: Rise of accuracy with multiple dynamic agents
Appendix C Additional Visualization of Results
In this section, to focus on the aggregation abilities of CAM, we have an oracle exploration where we explicitly say which rooms and items an agent has in its observations memory when answering the questions.
Below we provide additional visualizations of the results of our learning-based CAM methods against MV and debating on other Matterport environments. Like Table 1, we use independent answers of $3$ LLMs explored with LGX. Majority of the CAM methods consistently outperform the baseline majority vote (MV) and debate methods, with Decision Tree (DT) and XGBoost (XG) being the highest-performing methods.
Figure 3: 20 nodes
Figure 4: 43 nodes
Figure 5: 53 nodes
Figure 6: 78 nodes
Figure 7: 114 nodes
Figure 8: 145 nodes
Figure 9: Visualization of Decision Tree from Section 5.
Appendix D Decision Tree Visualization
In this section, we showcase the potential interpretability of our learning-based CAM framework by visualizing the decision tree trained in Section 5. This tree in Figure 9 was trained on queries only pertaining to the kitchen and was trained on independent answers from four different LLM agents. One agent had only observations from kitchen; another only dining room; another bedroom; and the last one only had office observations. We see that the main feature used was actually the object query. We provide the object encoding labels in Table 5.
Object Label Encoding $o$
appliance 0
armchair 1
bathtub 2
bed 3
board 4
bookcase 5
books 6
cabinet 7
chair 8
clothes 9
counter 10
curtain 11
cushion 12
desk 13
dresser 14
dumbells 15
hairbrush 16
headphones 17
jumprope 18
mirror 19
mug 20
nightstand 21
phone 22
picture 23
plant 24
playing cards 25
refrigerator 26
sink 27
sofa 28
table 29
television 30
toilet 31
toothbrush 32
towel 33
tv remote 34
wallet 35
water bottle 36
whiteboard 37
window 38
wristwatch 39
Table 5: Mapping of object labels to encodings $o$